### v4

# Kashmir-190006, India

Abstract
We consider scenarios where Jackiw-Teitelboim black holes are attached to baths whose temperatures can be manipulated externally. We consider the bath to be cooled continuously and numerically investigate the subsequent evolution of the generalized entropies of the black hole. Page curves corresponding to several different cooling profiles of the bath are obtained.
1 Introduction
The evaporation of black holes has been a topic of fascination for physicists for many years. According to the principle of unitarity, the entanglement entropy of radiation should initially increase but eventually decrease during the evaporation of a black hole, following what is now known as the Page curve Page:1993df. Obtaining a Page curve for black hole evaporation remained an open challenge for a long time. Only recently was it shown that for certain black holes, the entanglement entropy during evaporation indeed follows the Page curve Almheiri_2019; Penington:2019npb; Almheiri:2019hni.
In Almheiri:2019hni, for example, the set-up considered was a two-dimensional black hole in Jackiw-Teitelboim gravity attached to a bath. By computing the quantum extremal surfaces via the generalized entropy formula of Engelhardt and WallEngelhardt:2014gca, the Page curve was recovered. Since then, there has been significant progress in understanding the results and extending them to various black holes Penington:2019kki; Almheiri:2019qdq; Almheiri:2020cfm; Almheiri:2019yqk; Marolf:2020xie; Almheiri:2019psy; Chen:2020uac; Hashimoto:2020cas; Hartman:2020khs; Gautason:2020tmk; Balasubramanian:2020xqf; Krishnan:2020fer; Krishnan:2020oun; Ling:2020laa; Ahn:2021chg.
Most work on this topic has focused on black holes interacting with a bath that remains at a fixed temperature. However, the evaporation of a black hole can be influenced by external conditions such as changes in ambient temperature or energy exchange with other systems. For instance, primordial black holes in the early universe evaporate in an environment with rapidly changing ambient temperature. A natural question to ask is: how does changing the external conditions in a time-dependent manner affect the Page curve? In other words, how robust is the Page curve under time-dependent externalities?
In this paper, we address this question in the same setup considered in Almheiri:2019hni: a 2D black hole coupled to a reservoir. In our case, the reservoir is taken to be a cooling bath: one with controllable, time-dependent temperature. This provides a simple toy model for non-equilibrium evaporation.
The results from this model carry over beyond the case of a cooling bath. They apply to all processes where the black hole loses (or gains) mass in a time-dependent manner. Examples include the injection of negative mass shells into the black hole Cadoni:2021ypx and partially reflecting boundary conditions with time-dependent reflectivity. As we will show, all such setups for a black hole out of equilibrium with the bath are dependent on a single parameter, and hence mathematically equivalent.
In JT gravity, the black hole geometry is described by boundary parametrization $f(t)$. As the temperature of the bath changes, it leads to a change in $f(t)$. This in turn shifts the location of the quantum extremal surface (QES), thereby modifying the Page curve. This is the modification that we aim to study in this paper. There are a few special situations where analytic control is possible, such as instantaneous temperature jumps sourced by shock waves from the bath to the black hole Hollowood:2020cou. Outside of these cases, the evolution of $f(t)$, and therefore the Page curve, can not be computed analytically.
We therefore adopt a numerical approach that allows for arbitrary profiles beyond analytic solvability. We consider scenarios where the bath temperature is allowed to change continuously with time and compute the modified Page curves numerically for various cooling profiles. In the limiting case of shock waves, we demonstrate that our results are in agreement with those of Hollowood:2020cou, where they were computed analytically.
In more detail, we consider three different cooling profiles. They are: an instantaneous temperature drop (for the bath), linear drop in energy, and the temperature following Newton’s law of cooling. These profiles are chosen to model environments where a black hole exchanges energy with an external system whose properties evolve in a controllable way. Our only assumption is that the process is slow enough to remain within the semiclassical regime (back-reaction remains negligible).
For each case, we track the subsequent evolution of the boundary parametrization numerically utilizing the RK4 method, starting from the parametrization corresponding to a black hole at temperature $\beta$ and reaching a temperature $\tilde{\beta}$. We compute the positions of the quantum extremal surfaces and the corresponding entropies at each time step during the relaxation to a new temperature. The results are the Page curves of evaporating black holes in their non-equilibrium phases for each of the three cooling profiles.
Note that while we have restricted ourselves to these profiles, the numerical method we use can be utilized for any general case as long as it stays in the semiclassical regime.
To summarize, our main new contribution is extending the computation of the Page curve to cases where the black hole is interacting with a controlled, time-dependent environment (which changes the boundary parametrization in a continuous manner) utilizing numerical methods.
The structure and outline of the paper are set out as follows. In Section II, we start with preliminaries and provide the setup. Section III presents the computations for the instantaneous drop in temperature. In Section IV, we generalize to more general cooling profiles. We conclude with a brief summary of the new results and future directions.
2 Preliminaries
2.1 Jackiw-Teitelboim Gravity
We start by reviewing the basic aspects of Jackiw-Teitelboim(JT) gravity Jackiw:1984je; Teitelboim:1983ux and nearly AdS2 holography Maldacena:2016hyu; Almheiri:2014cka; Jensen:2016pah; Maldacena:2016upp; Engelsoy:2016xyb. We refer the reader to Mertens:2022irh for a recent detailed review.
The action of JT gravity minimally coupled to matter is given by:
$I[g,\Phi]=I_{\text{grav}}[g,\Phi]+I_{\text{matter}}$ (1)
Here, $\Phi$ denotes the dilaton field. The gravitational action is given by:
$I_{\text{grav}}[g,\Phi]=\frac{-1}{16\pi G_{N}}\int_{\mathcal{M}}\sqrt{g}$ $\Phi\left(R+2\right)-\frac{1}{8\pi G_{N}}\int_{\partial\mathcal{M}}\sqrt{h}\Phi(K-1)$
$-\frac{\Phi_{0}}{16G_{N}\pi}\left(\int_{\mathcal{M}}\sqrt{g}R+2\int_{\partial\mathcal{M}}\sqrt{h}K\right)$ (2)
The dilaton and the metric equations of motion are given by:
$R=-2,$ (3)
$\nabla_{\mu}\nabla_{\nu}\Phi-g_{\mu\nu}\nabla^{2}\Phi+g_{\mu\nu}\Phi=-8\pi G_{N}T_{\mu\nu}.$ (4)
where $T_{\mu\nu}$ is the stress-energy tensor of the matter sector. The dilaton equation of motion constrains the two-dimensional geometry to be locally AdS2.
2.2 Quantum Extremal Surfaces
The entropy of Hawking radiation and the generalized entropy are given, respectively, by the following formulae:
$S_{R}$ $=\min_{I}\text{Ext}_{I}\left[\frac{A(\partial I)}{4G_{N}}+S_{\text{QFT}}(R\cup I)\right],$ (5)
$S_{\text{gen}}$ $=\frac{A(\partial I)}{4G_{N}}+S_{\text{QFT}}(R\cup I).$ (6)
Here, $G_{N}$ denotes Newton’s constant, $R$ is the radiation region, $I$ is the island, $\partial I$ is the boundary of the island, and $S_{\text{QFT}}(R\cup I)$ represents the entanglement entropy of the quantum fields in the bulk region of spacetime. The term $R\cup I$ denotes the union of the island and the radiation region. Extremization of $S_{\text{gen}}$ determines the location of the island. The surfaces that extremize the generalized entropy are called quantum extremal surfaces (QES).
[…]
Figure 1: Quantum Extremal Surface
When there is a non-trivial QES in the bulk, it corresponds to an island. Figure 1 shows a QES where the exact position of ‘a’ is determined by extremizing the generalized entropy.
2.3 Black Hole at Equilibrium
We now introduce the setup of a black hole coupled to a bath. The framework of JT gravity black holes with leaky boundary conditions was originally developed in Engelsoy:2016xyb; Mertens:2019bvy.
In our case, we have the AdS2 Poincaré patch with coordinates $x^{\pm}$:
$ds^{2}=-\frac{4\,dx^{+}\,dx^{-}}{\left(x^{+}-x^{-}\right)^{2}}.$ (7)
where we have
$u=\frac{x^{+}+x^{-}}{2}\in\mathbb{R}\quad\text{and}\quad s=\frac{x^{+}-x^{-}}{2}>{\epsilon}$
The AdS2 boundary is parametrized by a coordinate transformation $u=f(t)$ that relates the Poincaré coordinates $x^{\pm}$ to the Rindler coordinates $y^{\pm}$. The metric in the $y^{\pm}$ coordinates is given by:
$ds^{2}=\frac{-4f^{\prime}(y^{+})f^{\prime}(y^{-})}{\left(f(y^{+})-f(y^{-})\right)^{2}}\,dy^{+}\,dy^{-}$ (8)
We attach a half-Minkowski thermal bath at the holographic time-like boundary at $z=-\epsilon$. Note that this serves as our IR cutoff. We use flat coordinates for the bath and we do so by analytically continuing the coordinates of the bulk.
$ds^{2}_{bath}=\frac{-1}{\epsilon^{2}}\,dy^{+}\,dy^{-}$ (9)
An eternal black hole with temperature $\frac{1}{\beta}$ has the boundary parametrization relating the Poincaré and black hole patches given by:
$x^{\pm}=\frac{\beta}{\pi}\tanh\left(\frac{\pi y^{\pm}}{\beta}\right).$ (10)
Then the black hole patch has a metric of the form:
$ds^{2}=-\frac{\pi^{2}}{\beta^{2}}\frac{4}{\sinh^{2}\left(\frac{\pi}{\beta}(y^{+}-y^{-})\right)}\,dy^{+}\,dy^{-}.$ (11)
The ADM energy of spacetime is holographically related to the boundary parametrization:
$E(t)=-\frac{\Phi_{r}}{8\pi G_{N}}\{f,t\},$ (12)
$\{f(t),t\}\equiv\frac{f^{\prime\prime\prime}(t)}{f^{\prime}(t)}-\frac{3}{2}\left(\frac{f^{\prime\prime}(t)}{f^{\prime}(t)}\right)^{2}.$ (13)
$\{f(t),t\}$ is the Schwarzian derivative of boundary parametrization. At the boundary, the rate of change of energy is:
$\frac{\partial E}{\partial t}=T_{y^{+}y^{+}}-T_{y^{-}y^{-}}$ (14)
Here $T_{y^{+}y^{+}}$ is the ingoing stress-energy tensor that originates from the bath attached to the black hole, and $T_{y^{-}y^{-}}$ is the outgoing stress-energy tensor, i.e., the Hawking radiation from the black hole.
Under a conformal transformation
$ds^{2}_{g}=\Omega^{-2}ds^{2}_{\eta}$
The stress-energy tensor components transform as follows:
$\left<T_{y^{\pm}y^{\pm}}\right>_{g}=\left<T_{y^{\pm}y^{\pm}}\right>_{\eta}-\frac{c}{12\pi}\frac{\partial_{\pm}^{2}\Omega}{\Omega},$ (15)
$\left(\frac{\partial x^{\pm}}{\partial y^{\pm}}\right)^{2}\left<T_{x^{\pm}x^{\pm}}\right>_{\eta}=\left<T_{y^{\pm}y^{\pm}}\right>_{\eta}+\frac{c}{24}\{x^{\pm},y^{\pm}\}.$ (16)
Equation (15) describes a Weyl transformation of stress-energy tensors, here $\left<T_{y^{\pm}y^{\pm}}\right>_{g}$ corresponds to the normal ordered stress tensor with respect to the frame with metric $ds^{2}=-\Omega^{2}(y^{+},y^{-})\,dy^{+}dy^{-}$, and $\left<T_{y^{\pm}y^{\pm}}\right>_{\eta}$ corresponds to the normal ordered stress tensor with respect to the frame with metric $ds^{2}=-dy^{+}dy^{-}$. Equation (16) represents a complete conformal transformation, where, $\left<T_{x^{\pm}x^{\pm}}\right>_{\eta}$ corresponds to the normal ordered stress tensor with respect to the flat metric $ds^{2}=-dx^{+}dx^{-}$, and $\left<T_{y^{\pm}y^{\pm}}\right>_{\eta}$ corresponds to the normal ordered stress tensor with respect to the flat metric $ds^{2}=-dy^{+}dy^{-}$ .
In the case of an eternal black hole in thermal equilibrium with the radiation bath, both the incoming and the outgoing modes of the stress tensor have temperature $\frac{1}{\beta}$
$\left<T_{y^{\pm}y^{\pm}}\right>=\frac{\pi c}{12\beta^{2}},$ (17)
Note that the expectation values from this point onward are normal ordered stress tensor with respect to a flat metric in the respective coordinate. We remove the subscript from the further discussion.
In equilibrium, we have
$\frac{\partial E}{\partial t}=0$ (18)
We can transform our stress tensor value using (16), and in doing so we get $\left<T_{x^{\pm}x^{\pm}}\right>=0$ as expected.
2.4 Entropy in terms of boundary parametrization
First, we introduce a new frame $w^{\pm}$ where the CFT is in a vacuum state with a flat background geometry:
$\left\langle T_{w^{\pm}w^{\pm}}\right\rangle=0$ (19)
In this frame, the matter contribution to entanglement entropy between two points $a$ and $b$ is given by the Cardy-Calabrese formula Calabrese:2004eu:
$S=\frac{c}{6}\log\left|\frac{(w^{+}_{a}-w^{+}_{b})(w^{-}_{a}-w^{-}_{b})}{\epsilon^{w^{+}_{a}}\epsilon^{w^{-}_{a}}\epsilon^{w^{+}_{b}}\epsilon^{w^{-}_{b}}}\right|.$ (20)
Here, $\epsilon^{w^{+}_{a}},\epsilon^{w^{-}_{a}},\epsilon^{w^{+}_{b}},\epsilon^{w^{-}_{b}}$} are the UV cutoffs for the entropy at points $a$ and $b$. The cutoffs are taken with respect to the proper distance in the physical metric.
Now we want to relate the $w\pmframe$ to the original coordinate system $y\pm$. The stress energy tensor in the two frames is related by:
$\left(\frac{dw^{\pm}}{dy^{\pm}}\right)^{2}\left\langle T_{w^{\pm}w^{\pm}}\right\rangle=\left\langle T_{y^{\pm}y^{\pm}}\right\rangle+\frac{c}{24\pi}\left\{w^{\pm},y^{\pm}\right\}$ (21)
First, consider the coordinate $w^{+}$.
We have from equations (19) and (21):
$\left\{w^{+},y^{+}\right\}=-\frac{24\pi}{c}\left\langle T_{y^{+}y^{+}}\right\rangle,$ (22)
Combining this with (38), we get:
$\left\{w^{+},y^{+}\right\}=-\frac{2\pi^{2}}{\beta^{2}},\quad y^{+}<0,$ (23)
Similarly, from equation (39):
$\left\{w^{+},y^{+}\right\}=-\frac{2\pi^{2}}{{\beta_{bath}}^{2}},\quad y^{+}\geqslant 0$ (25)
During $y^{+}<0$, $w^{+}=e^{\frac{2\pi}{\beta}y^{+}}$ solves (23) and provides a coordinate system for both the black hole and the bath. After $y^{+}=0$, the coordinate relation between $y^{+}$ and $w^{+}$ is to be obtained by solving (25).
As (25) is a third-order differential equation, we have to specify three initial conditions at $y^{+}=0$. We impose continuity in the $w^{+}$ coordinates up to the second-order derivative.
This choice comes from the fact that a discontinuity in $w,w^{\prime}$ or $w^{\prime\prime}$ would correspond to unphysical discontinuities in the geometry, while the change (may be discontinuous) in the third derivative is precisely what encodes the change (may be instantaneous) in energy flux from the bath.
This matches the physical expectation that a sudden temperature change should not introduce discontinuities in the geometry. Rather, it should produce a finite change in the stress tensor flux.
This gives us:
$w^{+}|_{y^{+}=0}$ $=1$ (26)
$\frac{dw^{+}}{dy^{+}}\Big|_{y^{+}=0}$ $=\frac{2\pi}{\beta}$
$\frac{d^{2}w^{+}}{dy^{+2}}\Big|_{y^{+}=0}$ $=\frac{4\pi^{2}}{\beta^{2}}$
For the case of discontinuous change in the bath temperature, the discontinuity is reflected as the discontinuity of the Schwarzian in the third-order derivative. This discontinuity is an artifact of our rather artificial assumption of a discontinuous change in the bath temperature.
A similar argument follows for $w^{-}$ coordinates, where we have
$\left\langle T_{y^{-}y^{-}}\right\rangle=\frac{\pi c}{12\beta^{2}},\quad y^{-}<0,$ $\left\langle T_{y^{-}y^{-}}\right\rangle=g(t)\geqslant 0$ (27)
Here, $g(t)$ is obtained by solving (14) for a specific bath cooling profile. It follows that for $y^{-}<0$, the relation between $y^{-}$ and $w^{-}$ has the same form as (23):
$\left\{w^{-},y^{-}\right\}=$ $-\frac{2\pi^{2}}{\beta^{2}},\quad y^{-}<0$ (28)
This is solved by $w^{-}=-e^{\frac{-2\pi}{\beta}}$.
For $y^{-}\geq 0$ the form of the relation is different from (25), however:
$\left\{w^{-},y^{-}\right\}=$ $g(y^{-})\geqslant 0$ (29)
Here too, we choose initial conditions that ensure continuity up to the second order derivative:
$w^{-}|_{y^{+}=0}$ $=-1$ (30)
$\frac{dw^{-}}{dy^{-}}\Big|_{y^{+}=0}$ $=\frac{2\pi}{\beta}$
$\frac{d^{2}w^{-}}{dy^{-2}}\Big|_{y^{+}=0}$ $=-\frac{4\pi^{2}}{\beta^{2}}$
Unlike in the previous case, here the Schwarzian is never discontinuous. This reflects the fact that although the bath temperature may change suddenly, the black hole temperature changes in a continuous manner.
The cut-off distances in (20) are taken in terms of the physical metric, which is $\epsilon^{y^{\pm}}$. In the w-frame, these are given by:
$\epsilon^{w^{+}}\epsilon^{w^{-}}$ $=-dw^{+}dw^{-}$ (31)
$-dw^{+}dw^{-}$ $=-dy^{+}dy^{-}\left(\frac{dy^{+}}{dw^{+}}\right)\left(\frac{dy^{-}}{dw^{-}}\right)$ (32)
$-dw^{+}dw^{-}$ $=-\Omega^{2}dy^{+}dy^{-}\left(\frac{dy^{+}}{dw^{+}}\right)\left(\frac{dy^{-}}{dw^{-}}\right)\frac{1}{\Omega^{2}}$ (33)
$-dw^{+}dw^{-}$ $=\epsilon^{y^{+}}\epsilon^{y^{-}}\left(\frac{dy^{+}}{dw^{+}}\right)\left(\frac{dy^{-}}{dw^{-}}\right)\frac{1}{\Omega^{2}}$ (34)
$\epsilon^{w^{+}}\epsilon^{w^{-}}$ $=\epsilon^{y^{+}}\epsilon^{y^{-}}\left(\frac{dy^{+}}{dw^{+}}\right)\left(\frac{dy^{-}}{dw^{-}}\right)\frac{1}{\Omega^{2}}$ (35)
The blackening factor $\Omega^{2}$ for the metric in the y-frame is given by:
$\Omega^{2}=\frac{-4f^{\prime}(y^{+})f^{\prime}(y^{-})}{\left(f(y^{+})-f(y^{-})\right)^{2}}$ (36)
Where, $f$ is the boundary parametrization function. Note that for the point $a$ in the bulk, $\Omega^{2}$ varies, but for the point $b$ in the bath, it is a constant.
Next, we fix the dilaton. In the semi-classical limit, the dilaton is obtained by solving the dilaton equation (4). We solve it in the $x^{\pm}$ frame and then relate the result to the $y^{\pm}$ frame using $x^{\pm}=f(y^{\pm})$.To do this, we first determine $\left\langle T_{x^{\pm}x^{\pm}}\right\rangle$. From (16), we see that $\left\langle T_{x^{+}x^{-}}\right\rangle$ and $\left\langle T_{x^{-}x^{-}}\right\rangle$ both vanish. The non-zero component $\left\langle T_{x^{+}x^{+}}\right\rangle$ is given by:
$T_{x^{+}x^{+}}=\frac{1}{f^{\prime}\left(y^{+}\right)^{2}}\left(\frac{\pi c}{12\beta^{2}}+\frac{c}{24\pi}\left\{f\left(y^{+}\right),y^{+}\right\}\right).$
From (41), this equals:
$T_{x^{+}x^{+}}=\frac{1}{f^{\prime}\left(y^{+}\right)^{2}}\partial_{y^{+}}E\left(y^{+}\right)=-\frac{\Phi_{r}}{8\pi G_{N}}\frac{1}{f^{\prime}\left(y^{+}\right)^{2}}\partial_{y^{+}}\left\{f\left(y^{+}\right),y^{+}\right\}$
Using $\left\{f^{-1}\left(x^{+}\right),x^{+}\right\}=-\left(f^{\prime}\left(y^{+}\right)\right)^{-2}\left\{f\left(y^{+}\right),y^{+}\right\}$, we get:
$T_{x^{+}x^{+}}=-\frac{\Phi_{r}}{8\pi G_{N}}\partial_{x^{+}}^{3}f^{\prime}\left(y^{+}\right)$
Now we follow the argument as in Hollowood:2020cou to obtain the expression for the dilaton:
$\Phi=\Phi_{0}+2\Phi_{r}\left(\frac{f^{\prime\prime}\left(y^{+}\right)}{2f^{\prime}\left(y^{+}\right)}+\frac{f^{\prime}\left(y^{+}\right)}{f(y^{-})-f\left(y^{+}\right)}\right)$
Now we are ready to calculate the entropy of the Hawking saddle and Island saddle.
3 Dynamical Setup and Numerical Framework
In this section, we will set up the non-equilibrium problem of a black hole interacting with a bath of changing temperature. We will introduce our numerical methods and use them to obtain the Page curve for the case of an instantaneous change in temperature. We will see that our numerical results match with the analytical results obtained by Hollowood:2020cou to a very high degree of accuracy. This serves as a check of our numerical methods.
3.1 Breaking equilibrium: time-dependent bath temperature
The equilibrium in (14) is broken when the incoming stress tensor bath is modified in any manner, leading to the following:
$\frac{\partial E}{\partial t}=T_{y^{+}y^{+}}^{\text{modified}}-T_{y^{-}y^{-}}$ (37)
The energy of the black hole can be described as a function of the boundary parametrization. By plugging the incoming stress tensor, this boundary parametrization can be computed, and hence the entire evolution of the system can be tracked. Keeping all their parameters constant in (37), the dynamics of the black hole is completely captured by the time dependence of the incoming component of the stress tensor.
This implies that all situations where the stress tensor is modified in a time dependent way are mathematically equivalent. We study a cooling bath, but the method applies equally to for instance a situation where negative energy radiation is emitted from the black hole to the bath. Another example would be partially absorbing boundary condition. Since the evolution depends on a single parameter $T_{y^{+}y^{+}}^{\text{modified}}$, all of these different cases can be captured by (37).
The change in the bulk geometry because of this is captured in the boundary parametrization in the semi-classical regime. Note that we assume that there is no back reaction on the geometry.
3.2 Instantaneous change in bath temperature: Boundary parametrization
We now consider the case of an instantaneous change in the bath temperature. We start with the bath and the black hole in equilibrium at inverse temperature $\beta$ and then make a sudden change at boundary time $t=0$. Before the change, at $t<0$ or alternatively $y^{+}<0$, we have:
$T_{y^{+}y^{+}}=\frac{\pi c}{12\beta^{2}}$ (38)
The temperature of the bath instantaneously changes to $\tilde{\beta}$ at the boundary time $t=0$. This change propagates to the bulk causally. After $y^{+}>0$, we get :
$T_{y^{+}y^{+}}=\frac{\pi c}{12\tilde{\beta}^{2}}$ (39)
Note that this setup is mathematically completely equivalent to the one studied in Hollowood:2020cou, where a positive energy shock wave was sent to the black hole from the bath and the temperature of the black hole shifts and settles down to a new temperature. As a consistency check on our numerical methods, we will show that our results for boundary parametrization match with the analytical results of Hollowood:2020cou.
We can express the outgoing stress-energy tensor $T_{y^{-}y^{-}}$ as a function of parametrization:
$T_{y^{-}y^{-}}=-\frac{c}{24\pi}\{f(y^{-}),y^{-}\}.$ (40)
Energy conservation at the holographic boundary given by (14):
$\frac{\partial E}{\partial t}=\frac{\pi c}{12\tilde{\beta}^{2}}+\frac{c}{24\pi}\{f(t),t\}$ (41)
Upon substituting the parametrization with E, we get:
$\frac{\partial E}{\partial t}=\frac{\pi c}{12\tilde{\beta}^{2}}-\frac{G_{N}c}{3\Phi_{r}}E$ (42)
$E=e^{-kt}\int\left[\frac{\pi c}{12\tilde{\beta^{2}}}e^{kt}dt\right]$ (43)
where $k=\frac{G_{N}c}{3\Phi_{r}}$.
We solve this differential equation with the boundary condition that at $t=0$ the ADM energy is $E=\frac{\pi\Phi_{r}}{4\beta^{2}G_{N}}$, which was the original energy of the black hole before any evaporation took place. The third order differential equation for boundary parametrization is obtained as:
$\{f,t\}=-2\pi^{2}[\tilde{\beta}^{-2}-(\tilde{\beta}^{-2}-\beta^{-2})e^{-kt}].$ (44)
At early times, the black hole was in equilibrium with the bath at temperature $T_{\beta}$, so:
$f(y^{\pm})=\frac{\beta}{\pi}\tanh\left(\frac{\pi}{\beta}y^{\pm}\right)$ (45)
$f^{\prime}(y^{\pm})=\text{sech}^{2}\left(\frac{\pi}{\beta}y^{\pm}\right)$ (46)
$f^{\prime\prime}(y^{\pm})=-\frac{2\pi}{\beta}\sinh\left(\frac{\pi}{\beta}y^{\pm}\right)\,\text{sech}^{3}\left(\frac{\pi}{\beta}y^{\pm}\right)$ (47)
At very late times, it once again settled to equilibrium at temperature $T_{\tilde{\beta}}$, so:
$f(y^{\pm})\approx\frac{\tilde{\beta}}{\pi}\tanh\left(\frac{\pi}{\tilde{\beta}}y^{\pm}\right)$ (48)
$f^{\prime}(y^{\pm})=\text{sech}^{2}\left(\frac{\pi}{\tilde{\beta}}y^{\pm}\right)$ (49)
$f^{\prime\prime}(y^{\pm})=-\frac{2\pi}{\tilde{\beta}}\sinh\left(\frac{\pi}{\tilde{\beta}}y^{\pm}\right)\text{sech}^{3}\left(\frac{\pi}{\tilde{\beta}}y^{\pm}\right)$ (50)
The initial conditions are chosen to ensure continuity of $f(t)$ up to the second derivative.
$f|_{t_{0}=0}$ $=0$ (51)
$\frac{df}{dt}\Big|_{t_{0}=0}$ $=1$
$\frac{d^{2}f}{dt}\Big|_{t_{0}=0}$ $=0$
Note that in general $\{f,t\}=g(t)$, where for the single instant drop in bath temperature gives the expression $g(t)=-2\pi^{2}\left[\tilde{\beta}^{-2}-\left(\tilde{\beta}^{-2}-\beta^{-2}\right)e^{-kt}\right]$. In general $g(t)$ would have to be solved from (14). The differential equation to solve for ‘f’ is the same as $w^{-}$(29), but with a slightly different initial condition as seen in (30) and (51). This makes them related by a Möbius transformation Hollowood:2020cou.
To ensure that the back-reaction on the geometry remains under control and that the semiclassical approximation is valid, we must have $k<<1$.
The exact solution of this differential equation was obtained in Almheiri:2019qdq; Hollowood:2020cou where $f(t)$ is expressible in terms of modified Bessel functions. The exact analytical solution for boundary parametrization is as follows:
$f(t)=\frac{\tilde{\beta}}{\pi}\frac{K_{\nu}(\nu z_{0})\left(\frac{\hat{f}(t)}{\hat{f}(t_{0})}-1\right)+z_{0}\tanh\frac{\pi t_{0}}{\tilde{\beta}}\left(\hat{f}(t)I^{\prime}_{\nu}(\nu z_{0})/\alpha-K^{\prime}_{\nu}(\nu z_{0})\right)}{K_{\nu}(\nu z_{0})\left(\frac{\hat{f}(t)}{\hat{f}(t_{0})}-1\right)\tanh\frac{\pi t_{0}}{\tilde{\beta}}+z_{0}\left(\hat{f}(t)I^{\prime}_{\nu}(\nu z_{0})/\alpha-K^{\prime}_{\nu}(\nu z_{0})\right)}$
where
$\hat{f}=\alpha\frac{K_{\nu}(\nu z)}{I_{\nu}(\nu z)},\qquad\nu=\frac{2\pi}{\tilde{\beta}k},\qquad z=\sqrt{\frac{E_{\text{shock}}}{E_{\beta}}}\,e^{-k(t-t_{0})/2}.$
The constant of normalization is fixed by using the initial condition on $\hat{f(t_{0})}=e^{2\pi t_{0}/\beta}$. Also for our case $t_{0}=0$. Here, the AdS radius is set to 1.
We now proceed to solve this equation numerically. The time step for evolution is taken to be 0.01. We use RK4 method for numerically solving the differential equation to obtain the boundary parametrization $f(t)$ and later for the $w\pm$’ coordinates.
First, we plot the first and second derivative of $f(t)$. This is shown in Fig 2(a) and Fig 2(b). As a consistency check, we match them with the analytically computed derivatives of $f(t)$ at early and late times.
(a) Plot of first derivative of $f(t)$ with $t$
(b) Plot of second derivative of $f(t)$ with $t$
Figure 2: Interpolation of derivatives
In the Figures 2(a) and 2(b), the numerically computed derivative is shown in red, while the analytically computed early and late time derivatives are shown in green and blue respectively. It can be seen how the derivative and double derivative of the boundary parametrization interpolate between the old and new boundary parametrizations.
(a) Plot of first derivative of $f(t)$ with $t$
(b) Plot of second derivative of $f(t)$ with $t$
Figure 3: Numerical check with analytical result
Figure 3(a) is a comparative plot between our numerical solution and the analytical solution of Hollowood:2020cou. The plots overlap entirely with only marginal accumulation error as shown in 3(b). This provides a check on our numerical methods.
RK4 typically has an error rate of the order $O(h^{4})$, where h (=0.01 in our case) is the time step. Our analysis is not very sensitive to this error; in fact, the standard Euler method for numerical integration also gives the same result. In order to satisfy the semiclassical regime, our k is set to “0.1” throughout the discussion. Other Parameters include $\beta=50\pi$ and $\tilde{\beta}=55\pi$, c=100 and $G_{n}=0.06$.
3.3 Instantaneous Change in Temperature: Entropy saddles
Now that we have computed the evolution of the boundary parametrization with time, we use it to compute the generalised entropy in the Hawking saddle and the Island saddle.
3.3.1 Hawking saddle
We start with the calculation of the Hawking saddle.
$b_{L}$$b_{R}$$b_{L}$$b_{R}$
Figure 4: Hawking saddle represented geometrically. Entropy is related to the proper distance between the two points.
Before calculating, we need to figure out the $w^{\pm}$ coordinates for both endpoints $b_{L}$ and $b_{R}$ (as shown in Figure 4). We start by specifying the points in the y frame and then write them in the w-frame coordinates. In the y-frame, we fix the spatial distance of the points from the cutoff surface. We fix $z=2\epsilon$, so at any boundary time t the $y^{\pm}$ coordinates of the points in the right bath are:
$y^{+}=(t+z)/2,\quad$ $y^{-}=(t-z)/2$ (52)
We proceed similarly for the left side bath and get these points in the corresponding $w^{\pm}$ coordinates. Note that the $w^{\pm}$ coordinates that we use to completely span the two-sided black hole, along with the two baths attached on either side. Then the Hawking saddle is the entanglement entropy between the two bath points using the Cardy-Calabrese formula (20).
Figure 5: Entropy of Hawking saddle vs Boundary time
In Figure 5, the entropy curves corresponding to the old and new black holes are the Hawking saddle entropy curve for eternal black holes corresponding to the initial and final temperature, respectively. It can be seen that the entropy curve for our evaporating black hole starts as the old one but later becomes parallel to the new saddle. As expected, the Hawking saddle entropy curve becomes linear at late times, with slope proportional to the temperature of the black hole. We have now obtained the graph for our evaporating black hole, which shows a clear interpolation between the initial and final states.
3.3.2 Island saddle
[…]
Figure 6: Geometric representation of island saddle. Entropy is related to the proper distance along the red lines and the value of dilaton at ‘a’. The points ‘a’ are QES in respect to the points ‘b’.
In the island saddle calculation, we iterate over all possible points ‘a’ ($y_{2}$’ in Figure 6) in a fixed time slice to find an extrema. Note that there is a symmetry between the left and right sides of the diagram.
Figure 7: Entropy of Island saddle vs Boundary time
Figure 7 shows the decrease in the island entropy over time from the black hole entropy corresponding to the black hole at the initial temperature to that corresponding to a black hole at the new temperature. The asymptotic behavior of the black hole coming to equilibrium in its final state is clearly reflected in the graph. As expected, the Page curve comes down for an evaporating black hole, whereas it stays at a constant value for an eternal black hole. Thus, we have directly obtained the Page curve after Page time for an evaporating black hole. In addition, this numerical approach can be used to profile evaporation for any type of cooling that the bath might provide.
4 Black Hole Evaporation: General Cooling Profiles
In this section, we explore different examples of cooling of the bath, which are continuous in nature and different from the earlier delta function drop in temperature.
The starting point for the general case is the generalization of (41):
$\frac{\partial E}{\partial t}=\frac{\pi c}{12\tilde{\beta_{t}}^{2}}+\frac{c}{24\pi}\{f(t),t\}$ (54)
where $\tilde{\beta_{t}}^{2}$ is the temperature of the bath.
We will focus on three examples: a case where the bath temperature follows Newton’s law of cooling, one where the incoming stress tensor from the bath has a drop in value linear in time, and finally, a quasi-static case.
4.1 Bath Energy follows Newton’s law of cooling
Newton’s law of cooling corresponds to the natural relaxation of a finite bath toward equilibrium with an external reservoir. It is widely used as an effective description in statistical mechanics.
In this case, we take $\frac{\pi c}{12\tilde{\beta_{t}}^{2}}=\frac{\pi c}{12}\left(\frac{1}{\tilde{\beta}^{2}}+\left(\frac{1}{\beta^{2}}-\frac{1}{\tilde{\beta}^{2}}\right)e^{-st}\right)$. We choose ‘$s$’ as the cooling constant. We first solve for ‘$E$’ and then solve for the boundary parametrization. The rest of the machinery works similarly, replacing $\tilde{\beta}$ with $\tilde{\beta_{t}}$.
We study how both the black hole entropy and the island entropy evolve for different relationships between the cooling constant $s$ and the parameter $k$. The three cases of interest are: $k>s,k<s$, and $k=s$. In each case, we also monitor the evolution of the black hole energy. For concreteness, we present numerical results for two representative choices: $s=k/2$ and $s=2k$.
(a) Black hole energy vs Time
(b) Island entropy vs Time
Figure 8: Energy and Entropy vs Time
In the figures 8(a) and 8(b) we see the change in energy of the black hole and its island saddle entropy, respectively. It appears that the Island entropy curve follows the energy curve. This is expected: the energy decreases, so does the area of the black hole. The island entropy decreases accordingly.
(a) Rate of change of energy
(b) Rate of change of island entropy
Figure 9: Rate of change of Energy and Entropy vs Time
Figure 10: Hawking Saddle Entropy
For Hawking radiation as well we can the slopes changing at different rates and finally converging to the slopes of the entropy curve for final state BH.
4.2 Linear Decrease in Incoming Stress Tensor
Linear decay of incoming stress tensor models scenarios in which the bath energy is drained at a constant rate. This provides a simple toy model that isolates the role of a time-dependent flux.
In this case, we take $\frac{\pi c}{12\tilde{\beta_{t}}^{2}}=\frac{\pi c}{12\beta^{2}}-st$. Here, $s$ is the rate of decay in the incoming energy. We follow the same procedure to calculate the island entropy for this evaporation profile.
Figures 8(a), 8(b), 10 show the change in energy, the Island saddle entropy, and the Hawking saddle entropy over time, respectively.
4.3 Quasi-static Evaporation
We would like to compare the results from the different models of evaporation with the quasi-static case. The quasi-static case is one where the equilibrium is maintained at each time step. We modeled this by treating the black hole as an eternal black hole at temperature $\beta_{t}$ at each time step, with $\beta_{t}$ taken as a constant during the timestep. The incoming and outgoing stress tensors are taken as the same function, thereby always ensuring equilibrium.
We can then vary $\beta_{t}$ as we like, and the entropy curve obtained will represent the island entropy for a quasi-static evolution. We take $\frac{1}{{\tilde{\beta}^{2}}_{t}}=\left(\frac{1}{\tilde{\beta}^{2}}+\left(\frac{1}{\beta^{2}}-\frac{1}{\tilde{\beta}^{2}}\right)e^{-kt}\right)$. This temperature profile corresponds to the temperature of a black hole evaporating due to a single sudden change in bath temperature, thereby breaking the equilibrium. If we use the same temperature profile for a black hole evaporating quasi-statically, the evolution corresponds to Newton’s law of cooling for both the bath and the black hole. Although we maintain the equilibrium with respect to 14, we further compute the boundary parametrization and the $w\pm$ coordinates numerically.
(a) Entropy of Island saddle vs Boundary time
(b) Entropy of Hawking saddle vs Boundary time
Figure 11: Quasi static vs Instantaneous (delta function shock) change in bath temperature.
5 Conclusion
In this paper, we extended previous results on Page curves for evaporating black holes in JT gravity to more general cases where the bath temperature can be controlled externally.
We computed the Hawking and the Island saddle entropy numerically for an evaporating black hole in an environment with varying temperature. Unlike previous discussions in which a black hole was given a shock and its evaporation back to its original state was studied, our approach is more general. We continuously change the bath temperature and consider several different profiles of continuous temperature change. In this paper, we studied three profiles: linear change, Newton’s law of cooling, and step function cooling. The result gives the Page curves of a slowly evaporating black hole in AdS under different external conditions of cooling, while remaining in the semiclassical regime.
Our numerical methods are entirely general and can be used to study any cooling profile of the bath. Physically, our result is expected to be useful for modeling black hole evaporation under general external conditions.
An interesting extension of our analysis would be to study stochastic fluctuations in the bath temperature, modeling random or noisy environments. Since our method only requires a time-dependent specification of the incoming stress tensor, incorporating stochasticity is straightforward and would allow one to probe how robust the Page curve is to environmental noise.
Another natural generalization concerns coupling the black hole to multiple baths with competing temperatures. This would allow for the exploration of steady-state non-equilibrium situations where radiation continuously flows through the black hole system.
We plan to return to these problems in the future.
6 Acknowledgments
NK was supported by the SERB Start-up Research Grant SRG/2022/000970 during a significant part of this research.